OpenStudy (anonymous):

can someone please check my answer What is the simplified form of x/2/4/2x i got 4x

OpenStudy (anonymous):

\[\frac{\frac{x}{2}} {\frac{4}{2x} } = \frac{x}{2}} \times {\frac{2x}{4} = \frac{x^2}{4}\] So \[\frac{x^2}{4} \] is the correct Ans.

OpenStudy (anonymous):

thanks i see how you did that